OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (precal):

part 2: y=2x-11 y=2x+3 same slope, lines will not cross, no solutions

OpenStudy (precal):

part 3: y=2x+4 y=(6/3)x+(8/2) same line, crosses in many places, many solutions

OpenStudy (anonymous):

thanks what about part 1?

OpenStudy (precal):

make it up, I made those up

OpenStudy (precal):

Just make sure they cross is one place

OpenStudy (anonymous):

can you do it please? :)

OpenStudy (precal):

y=2x+11 y=(-1/2)x+5 I made it perpendicular therefore they will cross in one place.

OpenStudy (anonymous):

thanks

OpenStudy (precal):

anytime :) don't be afraid to make some of them up yourself just check your solutions with a graphing calculator such a shame they wanted them in slope-intecept form

OpenStudy (anonymous):

i know right !! :)

OpenStudy (anonymous):

Is really helped me , thanks .